OpenStudy (anonymous):

i really need to know this please!!! Find the sum of the first 18 terms of the arithmetic sequence in which the first term is -9 and the 18th term is -60.

OpenStudy (anonymous):

\[a = -9\] \[a_{n} = a + (n-1)d\] a = first term n = number of terms d = difference between two terms \[a_{18} = -9 + (18-1) d\] \[-60 = -9 + (17) d\] Find d then sum of the series will be \[\sum_{k=0}^{n-1} (a+kd) = (n/2)(2a + (n-1)d)\]